### v1

# Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]

Abstract.
Graph Neural Networks (GNNs) have gained significant attention in recent years due to their ability to learn representations of graph-structured data. Two common methods for training GNNs are mini-batch training and full-graph training. Since these two methods require different training pipelines and systems optimizations, two separate classes of GNN training systems emerged, each tailored for one method. Works that introduce systems belonging to a particular category predominantly compare them with other systems within the same category, offering limited or no comparison with systems from the other category. Some prior work also justifies its focus on one specific training method by arguing that it achieves higher accuracy than the alternative. The literature, however, has incomplete and contradictory evidence in this regard.
In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems. We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations, with speedups between 2.4×2.4$\times$2.4 × - 15.2×15.2$\times$15.2 ×. We also find that both training techniques converge to similar accuracy values. Our work highlights the importance of comparing system across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for the comparison.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch.
1. Introduction
Graph neural networks (GNNs) are a class of machine learning models that reached state-of-the-art performance in many tasks related to the analysis of graph-structured data, including social network analysis, recommendations, and fraud detection (Zitnik et al., 2018; Schlichtkrull et al., 2018; Hamilton et al., 2018). They are often used to process large graphs that have millions of vertices and billions of edges (Ying et al., 2018; Hu et al., 2021; Liu et al., 2023). A large volume of recent work, both in academia and industry, has been dedicated to scaling GNN training to support such large graphs using multi-GPU systems. This is a challenging problem because GNNs run multiple rounds of message passing across neighboring vertices, which is an irregular computation.
Two classes of GNN systems: Full-graph and mini-batch. GNNs can be trained using either a mini-batch or a full-batch (typically called full-graph) approach, much like other machine learning models. In GNN training, however, these two approaches result in different message-passing patterns, which require different system pipelines and pose different system optimization challenges. This resulted in the development of two distinct classes of GNN training systems, each designed to support scalability for either mini-batch or full-graph training.
Figure 1. Different data management pipelines in two example systems: PipeGCN (full-graph) and DGL (mini-batch). The diagrams omit gradient synchronization.
The main reason for the emergence of these two separate class of systems is that different training methods (full-graph vs. mini-batch) require different data management pipelines to partition data, transfer it across CPUs and GPUs, and parallelize computation when scaling to large graphs. Full-graph GNN training performs message-passing for the entire batch, typically across the entire graph, at each epoch. To scale to large graphs that do not fit in the memory of a single GPU, multi-GPU full-graph training systems use model parallelism: they partition the graph, process different partitions at different GPUs, and perform message passing across partitions (Tripathy et al., 2020; Ma et al., 2019; Jia et al., 2020). For example, the PipeGCN (Wan et al., 2022b) full-graph training system partitions the input graph and keeps each partition in a different GPU (see Figure 1a). The black dotted lines show the communication of vertex features across GPUs during message passing at each GNN layer in the forward and backward pass. The communication happens through either a fast cross-GPU bus such as NVLink, if available, or PCIe.
Mini-batch training breaks up the training set into mini-batches. Each training epoch now consists of multiple iterations, one per mini-batch. Mini-batch training is popular in production systems because it is amenable to scaling through data parallelism (Zheng et al., 2020, 2022b; Zhu et al., 2019). A unique challenge when applying data parallelism to GNNs is that preparing mini-batches requires expensive pre-processing at each iteration. At each iteration, data-parallel mini-batch training systems sample the k-hop neighborhood of the vertices in the mini-batch, load each sample into a different GPU and perform local message passing only within each sample, without passing message across GPUs. For example, Figure  1b shows a training iteration in the popular DGL system. First, each GPU obtains a sample of the k-hop neighborhood of a subset of the vertices in the mini-batch. Sampling is preferably executed using the GPU since it is a computationally expensive step in itself (Jangda et al., 2021). Next, the sample and the input features of its vertices are loaded to the GPUs. Finally, each GPU performs a forward and backward pass on the local sample independently.
These different data management pipelines result in a different set of optimizations. Full-graph training systems must deal with expensive and irregular computation and high communication costs at each epoch. They propose sophisticated scheduling and partitioning algorithms to address these bottlenecks (Md et al., 2021; Peng et al., 2022; Wan et al., 2022b; Thorpe et al., 2021; Wan et al., 2023a). Work can be partitioned and scheduled offline, since the graph is static. Some full-graph training systems also propose techniques to reduce the volume of message data transferred across GPUs and hosts (Wan et al., 2023b; Peng et al., 2022; Wan et al., 2022a). In data-parallel mini-batch training systems, instead, computation at each iteration is much more lightweight and there is no cross-GPU message passing. However, pre-processing is a major bottleneck. Samples are typically large, so sampling them and loading them into GPUs is expensive. Research on scalable mini-batch systems has focused on addressing these bottlenecks by accelerating mini-batch sampling (Yang et al., 2022b; Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Cai et al., 2023; Zheng et al., 2022a, b) and reducing the cost of loading the input features by caching them in the memory of the GPUs (Lin et al., 2020; Yang et al., 2022b; Song et al., 2023b; Cai et al., 2023; Tan et al., 2023; Yang et al., 2022a; Liu et al., 2023; Polisetty et al., 2023).
Motivation: What is the state of the art? To understand the state of the art in the field it is necessary to compare systems across different classes. We performed an exhaustive review of the existing literature and found that prior works that propose systems in one class compare them with other systems in the same class, with limited or no performance evaluation of systems in the other class. Furthermore, several papers assert the superiority of the training approach they support in terms of the best accuracy that can be achieved, resulting in conflicting claims. We report a detailed discussion of the evidence in the literature in Section  2. Overall, the existing literature makes it difficult to get a clear picture of the state of the art of the field.
Performance evaluation. In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes (Wan et al., 2022b, a; Wang et al., 2019; Zheng et al., 2020; Tan et al., 2023). We chose systems that have a stable, publicly-available, and open-source implementation, are representative of the most important types of systems optimizations in their class, and implement all the standard GNN models (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018)) in their repository.
Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN model, and hardware deployments we considered. These results are consistent with the evidence in the literature. Several papers that propose full-graph training systems use epoch time as the main metric to compare with mini-batch training systems, showing that the latter have a larger epoch time  (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023). Our evaluation confirms these results. However, our results also highlight that considering the epoch time alone only gives a partial picture and that time-to-accuracy is a better metric for comparison. Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch. Therefore, even if mini-batch training systems perform more work per epoch than full-graph training ones to run multiple iterations, their overall time-to-accuracy is lower. Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider. Across all the datasets we considered, the fastest mini-batch training system is between 2.4×2.4$\times$2.4 × to 15.2×15.2$\times$15.2 × (6.3×6.3$\times$6.3 × on average) faster than the fastest full-graph systems in terms of time-to-accuracy.
In terms of scalability, full-graph training and mini-batch training systems observe similar relative speedups when they scale from a single GPU to multiple GPUs. However, the mini-batch systems we consider have a faster time-to-accuracy on a single GPU or host, so they remain faster in absolute terms in all configurations. We report our performance results in Section 4.
Performance vs. accuracy. The time-to-accuracy experiments discussed previously measure the time it takes different systems to reach the same target accuracy, which we select by performing hyperparameter tuning, using the same model and dataset. However, using a system with a longer time-to-accuracy can be justifiable if a model can converge to a higher accuracy. This is an important factor when navigating the performance trade-offs of using different GNN training systems. In fact, some papers proposing full-graph GNN systems assert that full-graph training achieves higher accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others proposing mini-batch training systems claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
Our results show that, with proper hyperparameter tuning, both full-graph and mini-batch training reach a similar accuracy across different datasets and models. Which method achieves the highest accuracy depends on the dataset and GNN model but the accuracy gaps are small. In our evaluation, we matched or exceeded all the accuracy results in the literature on GNN training systems we reviewed. We observe that given a dataset and GNN model, if we use the best hyperparameters for one training method with a system using the other training method, the latter achieves lower accuracy. This dependence on the choice of hyperparameters may explain the contradicting results found in the literature. Our results underline the importance of conducting separate hyperparameter tuning for the two training methods when comparing performance and accuracy across systems using different training methods. We also find that finding good hyperaparameters is easier with mini-batch training because it is less sensitive to variations in the model architecture. This has an impact on the overall training time, since model training often involves performing hyperparameter search. The harder it is to find good hyperparameters, the longer the overall training time. We report our accuracy experiments in Section 5.
Lessons learned. Our results show that the mini-batch training systems we consider consistently achieve a better performance than the full-graph training ones and a similar accuracy. We stress, however, that the goal of this study is not to establish the inherent superiority of one class of systems over the other. Our study is an empirical evaluation and the field is continuously evolving.
From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes. It also indicates a set of principles on how to perform this comparison: use time-to-accuracy as primary metric for comparison rather than epoch time, and perform separate hyperparameter tuning for mini-batch and full-graph training when selecting a target accuracy for performance comparison or comparing convergence accuracy. Prior works do not always follow these principles, and this may explain the apparent contradiction in the existing results. Our work offers guidance to ML practitioners who use GNN training systems and to systems researchers who design and evaluate them.
An extended version of this paper, which includes additional experimental results that we had to omit due to space constraints, is available at (Bajaj et al., 2024). To ease reproducibility, our study uses publicly available datasets, model implementations, and training systems. We make all our scripts publicly available. [^1] https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch
2. Background and Motivation
This section provides necessary background on GNN and GNN training systems to follow the rest of the paper. It then discusses the existing evidence in the literature.
2.1. Graph Neural Networks
Given a graph $G(V,E)$, where each node $v\in V$ is represented with a feature vector $h_{v}^{0}$ (input feature of vertex $v$) and the edges between any two vertices $u$ and $v$ is $e_{uv}\in E$. A GNN at $l^{th}$ layer performs the following computation to get the representation of any vector,
(1) $h_{v}^{l}=\Phi^{l}(h_{v}^{l-1},\Psi^{l}(h_{u}^{l-1},\forall u\in N\cup{v}),W^{l}),\$
where $h_{v}^{l}$ is the $l^{th}$ layer representation for vertex $v$, $N$ represents the incoming neighbor vertices of vertex $v$, $W^{l}$ is the model weight matrix, $\Psi^{l}$ is any aggregation function and $\Phi^{l}$ is an update function.
This paper considers the three standard GNN models used in previous work: GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2017). GraphSAGE learns the representation of a target node based on the features of the nodes in its k-hop neighborhood. The information from the neighboring nodes is propagated to the target node by performing message passing and aggregation recursively, one hop at a time. GAT (Graph Attention Networks) extend message passing by drawing inspiration from attention architectures in transformer models. It uses an attention score to weigh the incoming message of a node. GCN learns hidden layers representations and encode the graph structure and the features of the nodes by implementing a first-order approximation of spectral convolutions on graphs.
Table 1. Review of literature reporting claims of how mini-batch and full-graph training compare. (Notation: TTA - time to accuracy, ACC - accuracy, FG - Full graph, MB - mini-batch, ET - Epoch time)
Proposes System/paper Main Claims Experimental evidence Year Open Source Platform
Full-graph GNNPipe (Chen et al., 2023) FG lower ET ET 2023 Multi-host multi-GPU
Betty (Yang et al., 2023) FG higher ACC ACC 2023 Single-host single-GPU
HongTu (Wang et al., 2023) MB information loss, MB lower ACC ACC, TTA 2023 Single-host multi-GPU, distributed CPU-only
AdGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host single-GPU
G3 (Wan et al., 2023a) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
AdaQP (Wan et al., 2023b) MB information loss 2023 Mutli-host mult-GPU
NeutronStar (Wang et al., 2022) FG lower ET ET 2022 Multi-host single-GPU
PipeGCN (Wan et al., 2022b) FG higher ACC ACC 2022 Multi-host multi-GPU
BNS-GCN (Wan et al., 2022a) FG higher ACC, lower ET ACC, ET 2022 Multi-host multi-GPU
Sancus (Peng et al., 2022) MB information loss 2022 Multi-host multi-GPU
SAR (Mostafa, 2022) MB noisy gradients 2022 Multi-host CPU
Dorylus (Thorpe et al., 2021) FG higher ACC ACC 2021 Multi-host CPU, Multi-host single-GPU
DistGNN (Md et al., 2021) FG lower ET ET 2021 Multi-host CPU
DGCL (Cai et al., 2021) MB ACC loss 2021 Mutli-host mult-GPU
CAGNET (Tripathy et al., 2020) FG higher ACC 2020 Multi-host multi-GPU
ROC (Jia et al., 2020) FG lower TTA TTA 2020 Multi-host multi-GPU
NeuGraph (Ma et al., 2019) FG convergence guarantee 2019 Single-host multi-GPU
Mini-batch BGL (Liu et al., 2023) MB and FB same performance 2023 Multi-host CPU, Multi-host multi-GPU
NeutronOrch (Ai et al., 2023) FG impractical, limited GPU memory 2023 Single-host multi-GPU
DistDGLv2 (Zheng et al., 2022b) MB higher ACC, lower TTA ACC, TTA 2022 Multi-host CPU, Multi-host multi-GPU
SALIENT (Kaler et al., 2022) MB lower ET, higher ACC ACC, ET 2022 Multi-host multi-GPU
GNNLab (Yang et al., 2022b) FG hard to scale 2022 Single-host multi-GPU
ByteGNN (Zheng et al., 2022a) FG not practical for large graphs 2022 Multi-host CPU
CM-GCN (Zhao et al., 2021) MB lower TTA ACC, TTA 2021 Multi-host CPU
Blocking based (Yao and Li, 2021) FG higher memory and computational complexity 2021 -
DistDGL (Zheng et al., 2020) MB lower TTA 2020 Multi-host CPU, multi-host multi-GPU
PaGraph (Lin et al., 2020) MB and FB same performance 2020 Single-host multi-GPU
Cluster-GCN (Chiang et al., 2019) FG - memory: bad; time per epoch: good; convergence: bad 2019 -
LADIES (Zou et al., 2019) MB higher ACC ACC 2019 -
Other papers Rethinking (Li et al., 2024) FG not scalable ACC 2024 -
RDM (Kurt et al., 2023) FG ACC higher or equal, MB faster convergence ACC, TTA 2023 -
EXACT (Liu et al., 2021) FG - higher ACC for small graphs , MB - higher ACC for large graphs ACC 2021 -
OGB (Hu et al., 2020) MB higher ACC ACC 2020 -
2.2. Full-Graph vs. Mini-Batch Training: Different Systems for Different Pipelines
Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems. For a more detailed discussion on distributed GNN training systems, we refer the reader to (Shao et al., 2024; Lin et al., 2023).
Full-graph training systems The key idea behind full-graph training systems is to partition a large graph into multiple subgraphs such that each can fit into a GPU, and then train them in parallel. This requires communication of vertex features across GPUs. Different types of communication strategies have been used in the literature, including broadcast-based (Tripathy et al., 2020), point-to-point based (Md et al., 2021; Demirci et al., 2022), pipeline-based (Wan et al., 2023a; Mostafa, 2022; Md et al., 2021; Wang et al., 2021b), and communication via shared memory (Jia et al., 2020; Ma et al., 2019).
Communication incurs a substantial performance overhead and requires frequent synchronization. To overcome this bottleneck, two main types of optimizations have been proposed: asynchronous communication and sampling. Asynchronous communication allows the use of stale local information to avoid expensive synchronization costs. They can be categorized as epoch-fixed (Md et al., 2021; Wan et al., 2022b) and epoch-adaptive (Thorpe et al., 2021) staleness models. In GNN training, two synchronization points exist which are specific to GNN training: first, embedding aggregation and second, gradient aggregation. Some protocols (Peng et al., 2022; Md et al., 2021) only consider the embedding aggregation stage to support asynchronous aggregation, while others (Wan et al., 2022b; Thorpe et al., 2021) are designed to use historical information in both the embedding aggregation and gradient aggregation stages. Sampling keeps training synchronous but reduces communication cost by exchanging messages only among a subset of the boundary vertices across partitions (Wan et al., 2022a). Synchronous training is preferred in many cases because its results are predictable and reproducible.
Mini-batch training systems Mini-batch training systems rely on the generation of mini-batches to scale GNNs to large graphs and employ graph sampling to reduce the neighborhood explosion problem. Mini-batch training systems have a different set of performance bottlenecks than full-graph training ones. Existing research has focused on optimizing two steps, mini-batch sampling and input feature loading, which are major performance bottlenecks and often have a cost comparable to training. Multiple vertex-wise, layer-wise and subgraph-wise sampling techniques have been proposed (Chen et al., 2018a; Hamilton et al., 2017; Zeng et al., 2019; Zou et al., 2019). It is common to accelerate these algorithms using GPUs (Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Yang et al., 2022b). To avoid massive CPU-GPU communication while loading, some works (Zhu et al., 2019) have proposed caching techniques, like replicating frequently accessed remote vertices at local , store high out-degree vertices (Lin et al., 2020), caching hot vertices (i.e., those accessed more often) (Min et al., 2022). Some works (Liu et al., 2023) propose dynamic caching using a FIFO policy. Some distributed sampling techniques (Jiang and Rumi, 2021; Cai et al., 2023) prioritize local vertices, pushing sampling tasks to remote workers only if results need to be sent back, reducing communication. Some cache both topological and feature data on the GPUs (Sun et al., 2023; Zhang et al., 2023; Cai et al., 2023). To avoid graph sampling becoming a bottleneck, some works have focused on developing efficient GNN dataloaders (Bai et al., 2021; Deutsch et al., 2019; Dong et al., 2021; Min et al., 2021). Some systems reduce the cost of loading by proposing hybrid parallelism approaches for mini-batch training that are alternative to data parallelism: push-pull parallelism (Gandhi and Iyer, 2021) and split parallelism (Polisetty et al., 2023).
2.3. Motivation
The motivation for this work is that prior work on GNN training systems makes it difficult to establish the state of the art in the field. A significant amount of works on GNN training do not compare or evaluate their systems against systems using a different training approach(full-graph or mini-batch) (Kaler et al., 2023; Sun et al., 2023; Cai et al., 2023; Chen et al., 2018a; Hamilton et al., 2017; Chen et al., 2018b; Huang et al., 2018; Zeng et al., 2019; Gasteiger et al., 2022; Wang et al., 2021a; Zhu et al., 2019; Zhang et al., 2020; Kipf and Welling, 2017; Gandhi and Iyer, 2021). In the following, we review 33 papers on GNN training that do make claims comparing systems in the two classes. We summarize the claims in these papers in Table 1.
We classify the papers into three broad groups. The first group consists of papers that propose full-graph training. Out of the 17 papers in this group, 10 provide experimental evidence. The rest 7 papers make claims that full-graph training is superior in terms of higher accuracy, lower epoch time, and convergence guarantees and support the claims by only citing other papers. The second group includes 12 papers that propose mini-batch training. Among these papers, only 4 provide experimental evidence to support claims that full-graph training is difficult to scale and requires more memory, while mini-batch training has lower time-to-accuracy. The last group consists of 4 papers that do not propose one specific training approach. Overall, we find the following shortcomings in the existing literature, which motivate our comprehensive experimental comparison.
Conflicting claims about how the two approaches compare. Several papers proposing full-graph systems (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wang et al., 2023; Song et al., 2023a) demonstrate that their systems achieve higher accuracy and lower epoch time compared to mini-batch systems. Some papers mention that mini-batch causes information loss, noisy gradients, and does not guarantee convergence (Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019; Wang et al., 2023; Song et al., 2023a) However, papers proposing mini-batch systems (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020) and some other works (Hu et al., 2020) present contradicting evidence and claims showing that mini-batch training leads to higher accuracy and faster convergence to the target accuracy level. Other works claim full-graph training is impractical, requires high memory and is hard to scale (Ai et al., 2023; Yang et al., 2022b; Zheng et al., 2022a; Chiang et al., 2019; Yao and Li, 2021). Some argue that no approach is clearly superior and assessing their effectiveness requires a comprehensive evaluation which is out of scope of their work (Liu et al., 2021; Kurt et al., 2023; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Liu et al., 2023; Lin et al., 2020; Wang et al., 2023).
Comparing performance only in terms of epoch time. Certain works compare full-graph and mini-batch training systems based only on epoch time (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023). As we will show, a comparison in terms of time to accuracy is more informative given the differences in the two training approaches. Our observations indicate that although mini-batch training exhibits higher epoch time, it still takes less time to reach a target accuracy level (Section 4.1).
Experiment design. Some works (Jia et al., 2020; Liu et al., 2021; Li et al., 2024; Kaler et al., 2022) do not always use the same GNN model (e.g., GraphSAGE or GAT) for the two approaches as a basis of comparison. This conflates the effect of the training approach with the GNN model choice, since different models may exhibit varying performance even when trained with the same technique. Some papers do not mention explicitly whether they perform separate hyperparameter tuning for each method (Wan et al., 2023a, 2022b, 2022a; Thorpe et al., 2021; Zhao et al., 2021; Kaler et al., 2022). As we show in our experimental evaluation of Section 5, given a dataset and GNN model, it is possible to find hyperparameter settings where one method achieves higher accuracy than another and vice versa. Using hyperparameter tuning is necessary to perform a fair comparison.
Table 2. Dataset statistics. (*) The input features and the training data for Orkut are synthetic
Datasets #Nodes #Edges #Classes #Features Train/Val/Test
Pubmed 19,717 108,365 3 500 18,217/500/1,000
Ogbn-arxiv 169,343 1,166,243 40 128 90,941/29,799/48,603
Reddit 232, 965 11, 606, 919 41 602 152,410/23,699/55,334
Ogbn-products 2,449,029 61,859,140 47 100 195,922/48,980/2,204,127
Orkut 3,072,441 117,185,083 2* 128* 1,843,464/614,488/614489*
Ogbn-papers100M 111,059,956 1,615,685,872 172 128 1,207,179/125,265/214,338
Table 3. Hyperparameter search space
Parameter range/values
Number of Hidden Layers [2, 10]
Size of hidden layers [16, 1024]
Aggregator (GraphSage only) { mean, gcn, pool }
Number of heads (GAT only) [1, 12]
Sampling Fanout (MB only) [4, 50]
Learning rate [0.0001, 0.01]
Dropout [0.2, 0.9]
Mini-batch size (MB only) [128, 4096]
3. Experimental setup
Hardware setup. We run our experiments using two types of hosts. The first is a lower-end one, which we call PCIe. It has 2 Intel Xeon E5-2620 v3 CPUs with 12 cores each, 256GB of host memory, and 4 NVIDIA Tesla m40 GPUs, each having 24GB of memory, connected via PCIe. These servers are connected with a 1 GBps network. The second host type is higher end and we call it NVLink. It has 2 Intel Xeon Platinum 8480+ CPUs with 56 cores each, 256 GB of host memory, and 4 A100 GPUs, each with 80GB of memory. These servers are connected with a 25 GBps network. By default, all our experiments use 4 GPUs per host.
Datasets. Our experiments use six datasets: pubmed (Namata et al., 2012), ogbn-arxiv (Hu et al., 2020), reddit (Hamilton et al., 2017), ogbn-products (Hu et al., 2020), orkut (Yang and Leskovec, 2012) and ogbn-papers100m (Hu et al., 2020) (see Table 2). Except orkut, all these datasets come with input features that enable evaluating the accuracy of GNN training. The orkut dataset contains only the graph topology, so we only use it to measure epoch time for the scalability experiments. To run GNN training on it, we add synthetic input features of size 128, a train/val/test split of 60%/20%/20%. One out of two target classes is randomly assigned to each node.
Representative GNN training systems. To compare across full-graph and mini-batch distributed GNN systems, we select representative systems that incorporate the main optimizations discussed in Section 2.2. We chose systems that have a stable open-source implementation, including implementations of our target models (GCN, GraphSage, and GAT) in their repositories [^2] DGL 1.1.2: https://github.com/dmlc/dgl; Quiver 0.1.0: https://github.com/quiver-team/torch-quiver; DistDGL: https://github.com/dmlc/dgl/tree/master/examples/distributed; PipeGCN: https://github.com/GATECH-EIC/PipeGCN; BNS-GCN: https://github.com/GATECH-EIC/BNS-GCN. We run all systems on top of PyTorch 2.0.1 and Python 3.8.10.
For full-graph training, we select PipeGCN (Wan et al., 2022b), which supports asynchronous training to overlap communication and computation, and BNS-GCN (Wan et al., 2022a), which uses sampling to reduce communication. Both systems partition the entire graph topology and features, and store each partition in the memory of one GPU.
For mini-batch training, we choose popular data-parallel systems that support GPU-based sampling, efficient input feature loading, and feature caching. Specifically, we use DGL (Wang et al., 2019), which only supports single-GPU caching, and Quiver (Tan et al., 2023), which supports distributed caching. DistDGL (Zheng et al., 2020) is built on top of DGL and performs mini-batch training on a distributed cluster of hosts. It only supports CPU-based training and does not support caching. All the mini-batch systems we consider support only synchronous training.
Model selection and implementations. We consider standard GNN models that are commonly used to evaluate training systems: GCN, GraphSage, and GAT. We use the implementations provided with the distributions of the systems we benchmark and validate that the semantics of the implementations are consistent across different systems. This is to make sure that the model implementations are fully compliant and optimized for the respective systems and avoid introducing biases. DGL and Quiver include implementations of all the three models in its distribution. The PipeGCN distribution includes GraphSAGE, while the GCN and GAT implementations for both PipeGCN and BNS-GCN are included in the BNS-GCN distribution. For mini-batch training, we considered the standard Neighborhood Sampling algorithm for our evaluation (Hamilton et al., 2017).
Hyperparameter search space. The search space for hyperparameters is shown in Table 3. The first set of hyperparameters define the architecture of the GNN models. Fanout is related to the neighborhood sampling algorithm used in mini-batch training. Finally, we consider general model training parameters such as learning rate, dropout and batch size. To search this parameter space, we use random search for all the parameters to obtain the best set of hyperparameters giving the highest validation accuracy.
4. Performance Evaluation
We evaluate the performance of our representative GNN training systems and answer two questions: Q1) How long do representative systems take to reach the same target accuracy? Q2) How do these systems scale with a varying number of GPUs and hosts? We mainly use time-to-accuracy as the main metric for performance comparison since it is a more informative metric than epoch time, as we will discuss.
We conduct end-to-end performance measurements because mini-batch and full-graph training systems have different implementations and use different optimizations. For a more in-depth discussion of the effect of specific systems optimizations, it is necessary to restrict the evaluation to one class of systems, as done in (Yuan et al., 2024) for mini-batch systems.
4.1. Time-to-Accuracy
We start by answering question Q1: How long do representative systems take to reach the same target accuracy?
Methodology Time-to-accuracy is the time taken by any system to train to a target accuracy. To measure it, we select the target accuracies by leveraging hyperparameter tuning. Given a dataset and a GNN model, we first find a model architecture on which both training approaches (full-graph and mini-batch) converge to a similar and close-to-best accuracy, which we call convergence accuracy. We select the target accuracy as the minimum convergence accuracy across the two training approaches. We then use the same model architecture across all GNN training systems and measure their time to reach the target accuracy.
These experiments consider the two host types described in Section 3: a higher-end NVLink host and a lower-end PCIe host.
4.1.1. Single NVLink host
Table 4. Time to accuracy for GraphSAGE on a single NVLink host with 4 GPUs. Architectures (# layers, hidden size): pubmed - (3,256), ogbn-arxiv - (2,512), reddit - (3,1024), ogbn-products - (3,256), ogbn-papers100m - (2, 128). ET - Epoch time, TTA - Time to accuracy. For ogbn-papers100m DGL uses UVA mode for sampling and Quiver performs distributed caching. (Best results shown in bold.)
System pubmed ogbn-arxiv reddit ogbn-products ogbn-papers100m
Epoch time (s) PipeGCN (FG, Sync) 0.009 0.057 1.845 2.086 1.094
PipeGCN (FG, Async) 0.008 0.011 0.370 0.341 0.644
BNS-GCN (FG, Sync) 0.012 0.017 0.483 0.473 0.846
DGL (MB, Sync) 0.007 0.110 2.262 1.335 3.428
Quiver (MB, Sync) 0.005 0.131 1.994 2.117 1.878
Time to accuracy (s) PipeGCN (FG, Sync) 0.87 11.03 202.95 479.81 984.33
PipeGCN (FG, Async) 1.02 3.44 103.47 122.64 553.75
BNS-GCN (FG, Sync) 1.59 4.51 144.98 217.66 262.11
DGL (MB, Sync) 0.44 1.10 6.78 46.73 75.42
Quiver (MB, Sync) 0.36 0.65 9.97 84.70 41.31
Table 5. Time to accuracy with 3 NVLink hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached). (Best results shown in bold.)
System Time (s)
Epoch time (s) PipeGCN (FG, Sync) 1.139
PipeGCN (FG, Async) 0.352
BNS-GCN (FG, Sync) 0.432
DistDGL (MB, Sync) 4.452
Time to accuracy (s) PipeGCN (FG, Sync) 1093.63
PipeGCN (FG, Async) 306.59
BNS-GCN (FG, Sync) 133.92
DistDGL (MB, Sync) 35.62
Table 6. Time to accuracy and epoch time for GraphSAGE with a single PCIe host. Notation: ET - Epoch time, TTA - Time to accuracy. Architectures (# layers, hidden size): pubmed (3,256), ogbn-arxiv (2,512), reddit (4,1024) & ogbn-products (5,256). (Best results shown in bold.)
System pubmed ogbn-arxiv reddit ogbn-products
ET (s) PipeGCN (FG, Sync) 0.024 0.188 5.602 5.835
PipeGCN (FG, Async) 0.020 0.067 3.278 2.526
BNS-GCN (FG, Sync) 0.031 0.072 3.399 2.915
DGL (MB, Sync) 0.019 0.314 9.416 7.081
Quiver (MB, Sync) 0.019 0.216 10.272 9.246
TTA (s) PipeGCN (FG, Sync) 2.40 36.16 616.25 1341.95
PipeGCN (FG, Async) 2.35 20.75 917.80 909.43
BNS-GCN (FG, Sync) 2.91 19.33 1019.64 1340.80
DGL (MB, Sync) 1.25 3.14 28.25 247.83
Quiver (MB, Sync) 1.32 1.08 51.36 369.85
Table 7. Time to accuracy and epoch time with 3 PCIe hosts using the ogbn-papers100m graph and a GraphSAGE model with 2 layers and a hidden size of 128 (DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached). (Best results shown in bold.)
System Time (s)
Epoch time (s) PipeGCN (FG, Sync) 22.00
PipeGCN (FG, Async) 1.03
BNS-GCN (FG, Sync) 2.46
DistDGL (MB, Sync) 38.37
Time to accuracy (s) PipeGCN (FG, Sync) 19141.48
PipeGCN (FG, Async) 956.34
BNS-GCN (FG, Sync) 859.26
DistDGL (MB, Sync) 767.38
Figure 2. Time to accuracy for ogbn-arxiv, reddit, ogbn-products for GraphSAGE (single NVLink host).
Figure 3. Time to accuracy for ogbn-papers100M on (3 NVLink hosts)
Table 4 shows the time to accuracy when training GraphSAGE on a single NVLink host. Full-graph training systems have a consistently lower epoch time (ET) than mini-batch training systems because they perform only one training iteration per epoch. However, their time-to-accuracy (TTA) is substantially higher compared to mini-batch systems. For example, on the largest graph we consider, ogbn-papers100m, Quiver is 6.4$\times$ faster than BNS-GCN. Similarly, for pubmed, Quiver is 2.4×2.4$\times$2.4 × faster than PipeGCN sync (lowest speed-up we observed) and for reddit, DGL is 15.2×15.2$\times$15.2 × faster than PipeGCN async (highest speed-up we observed). On average, the fastest mini-batch systems are 6.3×6.3$\times$6.3 × faster than the fastest full-graph system we considered for these datasets. This is because mini-batch training systems perform multiple training iterations and model updates per epoch and thus converge in fewer epochs.
Figure 2 shows the time-to-accuracy curves. The dotted line indicates the target accuracy. We observe mini-batch training consistently has a much faster time to accuracy across graphs with different sizes. Results on other models are shown in the extended version (Bajaj et al., 2024).
The experiments also show the impact of the systems optimizations introduced in previous work. Among the full-graph training systems, PipeGCN can substantially decrease epoch time and time-to-accuracy using asynchronous training compared to its baseline synchronous version. Sampling, as proposed by BNS-GCN, is less effective than asynchronous training on the smaller graphs, but it yields the fastest time-to-accuracy among all full-graph training systems on ogbn-papers100m.
All full-graph systems we consider fully cache the graph in GPU memory. Mini-batch training systems also cache all the graph in the GPU memory with the exception of ogbn-papers100m, which cannot be cached by DGL. This is because DGL only supports caching when the entire graph fits in the memory of each GPU, unlike Quiver which implements a distributed cache where the graph can be partitioned among multiple GPUs. Despite the lack of caching, DGL still achieves a lower time-to-accuracy than the full-graph training systems on ogbn-paper100m.
4.1.2. Multiple NVLink hosts
We perform a similar experiment using three NVLink hosts with the largest of the graphs we consider, ogbn-papers100m.
Table 5 shows the results. Figure 3 shows the corresponding time-to-accuracy curves. Quiver does not support multi-node training, so we do not include it in these experiments. We also replace DGL with its distributed version, DistDGL. The main trends are similar to the ones observed for the single-host experiments. DistDGL, which uses mini-batch training, has a much larger epoch time than the other systems, which use full-batch training. Its time-to-accuracy, however, is much shorter than the other systems. It is worth noting that DistDGL has a much larger time-to-accuracy than DGL on a single host (see Table 4). This is because DistDGL uses CPU-based sampling and does not support caching, unlike DGL.
4.1.3. PCIe hosts
Table 6 reports the results for single-host training with GraphSAGE using a lower-end host setup, where NVLink is not available and cross-GPU communication within one host must occur over a slower PCIe bus. Compared to the NVLink experiments of Table 4, the epoch time of all systems is larger because the GPUs are less powerful. Full-graph training systems experience a stronger degradation in performance than DGL because they must run cross-GPU communication over PCIe, whereas DGL avoids communication by using data parallelism. Quiver also observes a performance degradation since loading data from other GPU’s cache is not as effective over the PCIe bus as it is over NVLink.
Table 7 futher compares all systems using 3 PCIe hosts for the ogbn-papers100m graph. In the PCIe setup, full-graph training systems need a multi-node setup to fully cache the ogbn-papers100m graph, which is larger than the others. In the setup with 3 PCIe hosts, hosts are connected by a slower network (1 GBps instead of 25 GBps). Like in the previous experiments, DistDGL has a lower time-to-accuracy than the other full-graph training systems, but the difference is much smaller than with the NVLink setup. This is because DistDGL requires cross-host communication to perform sampling and feature loading. These costs become bottlenecks over the slow 1GBps network of the PCIe host. By contrast, full-graph training systems are optimized to minimize communication and amortize its cost.
Takeaway. Mini-batch training enables faster time-to-accuracy compared to full-graph training since it requires fewer epochs to reach a target accuracy. This is because full-graph training systems update the model only once per epoch, whereas mini-batch training systems perform multiple iterations per epoch and update the model at each iteration. By updating the model more frequently, mini-batch training systems typically converge faster than full-graph training ones, even if their epoch time is longer.
These findings stress that time-to-accuracy measurements should always be included to provide a comprehensive comparison of systems using different training approaches. Comparing only in terms of epoch time, as is commonly done in prior work (see Table 1) does not provide a comprehensive comparison.
4.2. Scalability
Table 8. Single-host scalability with the GAT model using the NVLink host. Times in seconds. (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 6, 25), Reddit - (2, 1024, 4, 20), ogbn-products - (3, 128, 3, 20), Orkut - (3, 128, 2, 5). ET - Epoch time, TTA - Time to Accuracy.
Dataset ogbn-arxiv reddit ogbn-products orkut
System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET ET speed-up
PipeGCN (Sync) 1 1.293 569.01 1.0x 4.265 4094.59 1.0x 2.058 823.36 1.0x 1.898 1.0x
2 0.868 381.74 1.5x 2.566 2463.65 1.7x 1.341 536.20 1.5x 1.873 1.0x
3 0.763 335.90 1.7x 2.587 2483.90 1.6x 1.129 451.64 1.8x 1.594 1.2x
4 0.625 275.04 2.1x 1.801 1728.58 2.4x 1.033 413.28 2.0x 1.442 1.3x
PipeGCN (Async) 1 1.292 581.27 1.0x 4.268 4267.70 1.0x 2.060 782.76 1.0x 1.898 1.0x
2 0.777 349.83 1.7x 2.309 2309.20 1.8x 1.135 431.34 1.8x 1.149 1.7x
3 0.589 265.10 2.2x 2.309 2309.20 1.8x 0.810 307.61 2.5x 0.920 2.1x
4 0.469 211.14 2.8x 1.312 1312.15 3.3x 0.614 233.24 3.4x 0.720 2.6x
BNS-GCN (Sync) 1 1.309 471.06 1.0x 4.189 4188.60 1.0x 2.171 455.95 1.0x 2.173 1.0x
2 0.764 274.93 1.7x 2.315 2315.30 1.8x 1.213 254.73 1.8x 1.284 1.7x
3 0.532 191.56 2.5x 1.724 1723.60 2.4x 0.857 179.99 2.5x 0.984 2.2x
4 0.420 151.20 3.1x 1.395 1395.40 3.0x 0.629 132.13 3.5x 0.713 3.0x
DGL (Sync) 1 15.006 150.06 1.0x 24.006 360.09 1.0x 10.554 158.31 1.0x 14.791 1.0x
2 7.290 72.90 2.1x 12.794 191.91 1.9x 5.827 87.40 1.8x 7.668 1.9x
3 5.550 61.04 2.5x 8.949 134.24 2.7x 3.894 58.42 2.7x 4.940 3.0x
4 3.871 42.58 3.5x 6.510 97.65 3.7x 2.730 40.95 3.9x 3.980 3.7x
Quiver (Sync) 1 13.813 138.13 1.0x 26.427 396.41 1.0x 11.212 168.19 1.0x 13.718 1.0x
2 6.575 65.75 2.1x 13.764 206.46 1.9x 5.797 86.95 1.9x 7.447 1.8x
3 4.979 49.79 2.8x 8.703 130.55 3.0x 3.993 59.90 2.8x 5.058 2.7x
4 3.596 35.96 3.8x 6.779 101.68 3.9x 2.974 44.61 3.8x 3.855 3.6x
Table 9. Distributed scalability with the GraphSage model using 4 GPUs per host connected with NVLink . Times in seconds. (#Layers, hidden size, fanout): Orkut - (3, 512, 10), ogbn-products100M - (2, 512, 10). ET - Epoch time, TTA - Time to Accuracy. DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached
orkut ogbn-papers100m
System # Hosts ET ET Speed-up ET TTA TTA Speed-up
PipeGCN 1 4.769 1.0x 3.556 924.46 1.0x
(Sync) 2 4.588 1.0x 3.566 927.13 1.0x
3 4.351 1.0x 3.933 1022.63 0.9x
PipeGCN 1 0.454 1.0x 0.759 197.57 1.0x
(Async) 2 0.341 1.3x 0.464 120.67 1.6x
3 0.227 2.0x 0.361 93.96 2.1x
BNS-GCN 1 0.810 1.0x 0.997 209.38 1.0x
(Sync) 2 0.608 1.3x 0.762 160.10 1.3x
3 0.405 2.0x 0.542 113.88 1.8x
DistDGL 1 9.826 1.0x 11.561 80.93 1.0x
(Sync) 2 7.305 1.3x 10.749 75.24 1.1x
3 6.815 2.0x 6.612 46.28 1.7x
Next we answer Q2: How do systems scale with a varying number of GPUs and a varying number of hosts? For each dataset, we pick a GNN architecture that fits in a single GPU and use the single-GPU training performance as a basis to calculate the scalability speedups.
4.2.1. Single-host scalability (NVLink)
We now report the single-host scalability of our two approaches by varying the number of GPUs. We consider the GAT model since it is the most computationally demanding among the ones we considered.
Table 8 presents the results from our scalability experiment. The performance trends when training on a single GPU are similar to the ones we observed for the multi-GPU setups we discussed previously. The full-graph training systems have a much lower epoch time than the mini-batch ones but a higher time-to-accuracy.
When scaling to multiple GPUs, one important advantage of mini-batch training systems is that they can keep the amount of work per GPU constant by increasing the mini-batch size. This is not possible in full-graph training systems since the total amount of work per epoch is constant. As the number of GPU increases, the amount of work per GPU decreases but the communication cost increases, making it difficult to achieve scalability.
This effect is clear when comparing systems using synchronous training. With PipeGCN, which uses full-graph training, the cost of synchronous cross-GPU communication during the forward and backward pass of training hides the gains of reducing the computation cost per GPU.
Several optimizations in full-graph training systems address this problem. PipeGCN has a much better scalability when using asynchronous training. This enables PipeGCN to use pipelining, one of its most important features, to overlap communication and computation and to scale to up to 3.4×3.4$\times$3.4 ×. Sampling of boundary nodes in BNS-GCN also enables good scaling, up to 3.5×3.5$\times$3.5 ×, since the cross-GPU communication is avoided. The mini-batch systems scale almost linearly despite using synchronous training: both DGL and Quiver scale by up to 3.9×3.9$\times$3.9 ×. Besides scalability, these systems achieve lower time-to-accuracy that the full-graph training systems in all settings. We only show ET and ET Speed-up for orkut dataset, since it is a SNAP dataset (Yang and Leskovec, 2012) with only graph structure and synthetic input features and target labels.
We also experiment with GraphSAGE model, and the results are shown in the extended version (Bajaj et al., 2024). For GraphSAGE we see that synchronous PipeGCN doesn’t scale at all, asynchronous PipeGCN scales up to $\times3$ and BNS-GCN scales up to 2.7×2.7$\times$2.7 ×. While DGL and Quiver are able to scale up to 4.0$\approx$ 4.0$\times$ 4.0 × and 3.8×3.8$\times$3.8 × respectively.
Takeaway. The mini-batch training systems show better scalability than the full-graph training ones. This is because they can tune the mini-batch size to keep the work per GPU constant to maintain high GPU utilization, unlike the full-graph training systems. The full-graph training systems also experience increasing communication costs as the number of GPUs increases. However, these overheads can be mitigated using asynchronous training and boundary nodes sampling.
4.2.2. Distributed scalability (NVLink)
We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has four GPUs. We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution. We do not consider Quiver because it does not support distributed training.
Table 9 shows the results for distributed scalability with NVLink enabled devices. The synchronous version of PipeGCN cannot scale when increasing the number of hosts. Using asynchronous training and boundary nodes sampling can make full-graph training systems scale much better, as we also observed for the single-host experiments. DistDGL has similar scalability trends despite the fact that it does not use GPU caching. Despite its higher epoch time compared to full-graph training systems, DistDGL achieves a much lower time-to-accuracy.
Takeaway. The general scalability considerations we made for single-host scalability apply also to multi-host scalability. Synchronous full-graph training does not scale to multiple host. By using asynchronous training and boundary nodes sampling, PipeGCN and BNS-GCN achieve similar scalability as DistDGL, which uses mini-batch training. However, DistDGL has a lower time-to-accuracy in all configurations.
4.2.3. Single-host scalability (PCIe hosts)
We run the same experiments on PCIe hosts (as done in Section 4.1.3), and observe similar results, which are shown in the. We scale the number of GPUs from one to four.
All the full-graph training systems have a lower scalability because they require communication over the PCIe bus, which is slower than NVLink. This slowdown compared to the NVLink case is particularly evident for PipeGCN sync, which can only scale up to 2.0×2.0$\times$2.0 ×. PipeGCN async can scale up to at most 3.1×3.1$\times$3.1 × and BNS-GCN up to at most 3.0×3.0$\times$3.0 ×.
Mini-batch training systems have a better scalability because they require little cross-GPU communication. DGL scales up to 3.8×3.8$\times$3.8 × and Quiver up to 3.5×3.5$\times$3.5 ×.
Takeaway. The full-graph systems are more sensitive to the performance of the cross-GPU interconnect because they need to exchange vertex features between partitions at each GNN layer during the forward and backward pass. When GPUs must communicate over a slow PCIe interconnect, the scalability of these systems is lower. The mini-batch training systems use data-parallelism, which does not require exchanging vertex features across GPUs. Therefore, these systems show better scalability in the PCIe host.
5. Accuracy Evaluation
Even though a system takes longer than another to reach the same target accuracy, it could still be preferable to use it if it can converge to a higher accuracy. Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
Figure 4 summarizes the accuracy values reported in the literature we reviewed for different GNN models and datasets. Each point in the graph represents the test accuracy reported for a specific GNN model-dataset pair in one of the papers we reviewed. The figure illustrates a wide distribution of accuracy values for the same dataset and model combination.
In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the best accuracy we can achieve using full-graph and mini-batch training? Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method? In practice, the cost of training a model also depends on how long it takes to find good hyperparameters. Therefore, we also consider the following question: Q5) How sensitive are the two training methods to changes in the hyperparameters?
Figure 4. Test accuracy reported in literature on GNN training. (Wan et al., 2022b; dgl, 2023; Hu et al., 2020; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Md et al., 2021; Yang et al., 2023; Li et al., 2024; Gasteiger et al., 2022; Mostafa, 2022; Kaler et al., 2022; Yao and Li, 2021; Zou et al., 2019; Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2019; Chen et al., 2018b, a; Yao and Li, 2021; Wu et al., 2024; Veličković et al., 2018; Thorpe et al., 2021; Frasca et al., 2020; Wan et al., 2022a; Kipf and Welling, 2017; Jia et al., 2020)
Methodology To compare maximum accuracies among full-graph and mini-batch training, we use the methodology depicted in Figure 5. Given a dataset and a GNN model, we conduct two separate hyperparameter searches, one using full-graph training and one using mini-batch training. This is because the two training methods can have different sets of best hyperparameters for the same dataset and model. The space of hyperparameters we consider is shown in Table 3. The best hyperparameter settings we obtained using the two training methods are denoted as FG-search and MB-search. We then run full-graph and mini-batch training on both settings, measure the test accuracy, and denote it as FG-train and MB-train respectively. This yields a total of four accuracy combinations per dataset and GNN model. We report all the hyperparameter settings we found in the extended version of this paper (Bajaj et al., 2024).
To factor out any accuracy bias due to systems-level optimizations, we perform synchronous training with PipeGCN and DGL. We validate our results by verifying that given a dataset, model, and training method, our hyperparameter search achieves accuracy values that are equal to or better than the best accuracy value reported in the literature reviewed in Section 2.3.
Figure 5. Methodology of our accuracy comparison between full-graph (FG) and mini-batch (MB) training.
5.1. Accuracy with the Best Setting
We start by answering question Q3) What is the best accuracy we can achieve using full-graph and mini-batch training? Answering this question shows whether there is one method that consistently reach higher accuracy than another. When the hyperparameters are properly set, whether full-graph training achieves higher accuracy than mini-batch training or vice versa depends on the dataset-GNN model pairs.
The accuracy results for the GraphSAGE model are shown in Table 10. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
The results show that the accuracies are similar across training methods. For the GraphSAGE model, each method can yield higher accuracy than the other, depending on the dataset. The largest gap between the accuracies obtained with the two methods is less than $0.9\%$ in favor of full-graph training and averaging around $0.3\%$. We ran a similar analysis for the GCN and GAT model, and we report the results in the extended version of this paper (Bajaj et al., 2024). With the GAT model, the largest gap is slightly higher, up to $1.1\%$ in favor of full-graph training and the average is around $0.6\%$. Gaps are higher for GCN, which is a simpler model and turns out to be more sensitive to the choice of hyperparameters. The largest gap in this case is $3.6\%$ in favor of mini-batch training and the average is around $1.27\%$.
Takeaway. These results indicate that no training method alone guarantees higher accuracy than the other and the gap between the two methods is typically small (i.e., less than 3.6% accuracy). If achieving the highest accuracy is an absolute priority regardless of the training cost, it is advisable to conduct hyperparameter search twice using both training methods. Otherwise, using either of the two methods is likely to produce similar accuracy in the absence of other systems-level optimizations such as asynchronous training or boundary node sampling.
Table 10. Accuracy (%) of GraphSage using hyperparameter tuning method in Figure 5.
Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
FG-train $\textbf{77.40}\pm\textbf{00.04}$ $76.10\pm 00.67$ $\textbf{71.76}\pm\textbf{00.16}$ $67.21\pm 00.9$ $\textbf{96.94}\pm\textbf{00.01}$ $90.28\pm 00.63$ $\textbf{79.29}\pm\textbf{00.15}$ $76.60\pm 00.32$ $\textbf{62.82}\pm\textbf{00.01}$ $61.74\pm 00.13$
MB-train $77.00\pm 00.39$ $\textbf{78.20}\pm\textbf{00.32}$ $70.73\pm 00.38$ $\textbf{71.67}\pm\textbf{00.30}$ $96.52\pm 00.14$ $\textbf{96.82}\pm\textbf{00.04}$ $78.72\pm 00.07$ $\textbf{78.84}\pm\textbf{00.12}$ $62.52\pm 00.04$ $\textbf{62.77}\pm\textbf{00.02}$
5.2. Accuracy with Sub-Optimal Setting
We now answer the next question Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method? This shows how the outcome of the comparison between the two training approaches can change when using sub-optimal hyperparameter settings.
To answer this question for the case of the GraphSAGE model, we still look at Table 10. Given a dataset, the FG-search column reports the accuracy achieved with the two training methods using the hyperparameter setting where full-graph training works best. Mini-batch training performs worse than full-graph training, but only by a small margin. The largest gap between MB-train and FG-train is for the ogbn-arxiv graph ($1.03\%$).
The MB-search column compares the two training methods using the best hyperparameters we found for mini-batch training. Full-graph training has lower accuracy than mini-batch training and the gap between the two methods is larger, reaching $6.22\%$ for the reddit dataset.
Our evaluation on GAT and GCN, which is reported on the extended version (Bajaj et al., 2024), leads to similar conclusions. For GAT, largest difference between FG-train and MB-train is $2.85\%$ in favor of MB-search for the ogbn-arxiv dataset and $1.9\%$ in favor of FG-search for the pubmed dataset. For GCN, it is $1.36\%$ in favor of FG-search for the reddit dataset and $5.5\%$ in favor of MB-search for the pubmed dataset.
Takeaway. These results show that given a dataset and GNN model, each training method can outperform the other depending on the choice of the hyperparameters. This corroborates our choice of performing two separate hyperparameter searches, one per training method, for our comparison. It also potentially explains why existing work shows inconsistent evidence about which training method leads to higher accuracy. We hypothesize that the large variations in accuracy shown in Figure 4 is due to the difference in the choice of hyperparameter settings, which can greatly impact the accuracy and performance of GNN training.
Mini-batch training still performs relatively well in settings where full-graph training performs best, but the opposite is not true. This is an indication that mini-batch training is more robust to different hyperparameter settings, potentially making hyperparameter search easier. We will further corroborate this thesis in later experiments.
5.3. Ease of Hyperparameter Search
Figure 6. Sensitivity to the number of layers and hidden size for GraphSAGE model.
We finally answer our last question focusing on accuracy and hyperparameter search, which is Q5) How sensitive are the two training methods to changes in the hyperparameters? We evaluate the ease of finding a good hyperparameter setting using the two training approaches by measuring how accuracy varies as we vary the model architecture hyperparameters, i.e., the number of hidden layers and the size of the hidden vertex dimensions.
We trained each model for a maximum of 1000 epochs, stopping earlier if we the validation accuracy doesn’t increase for 50 epochs.
Figure 6 illustrates the best test accuracy when training the GraphSAGE model, top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training. Notably, mini-batch training demonstrates remarkable consistency across varying model architectures. In contrast, full-graph training exhibits a wider range of variations in test accuracy when the architecture changes. We observed similar results for GCN and GAT models, which are shown in the extended version of this paper (Bajaj et al., 2024).
Figure 6 and the other results for GCN and GAT, which are reported in the extended version (Bajaj et al., 2024), clearly show the effect of over-smoothing, which appears in GNNs at increased depths (Rusch et al., 2023). Mini-batch training on GraphSAGE does not show smoothing with up to 8 layers in Figure 6. In a set of separate experiments we found that oversmoothing does occur with mini-batch training, but only after we increase the number of layers to $16$ or higher. Our experiments for GCN and GAT show that oversmoothing occurs at lower depths for mini-batch training but the resulting accuracy drop is still much less than that with full-graph training (see the extended version (Bajaj et al., 2024)).
Takeaway. These results highlight the robustness of mini-batch training to alterations in model architecture. Full-graph training appears to be more sensitive to changes in architecture, so it requires a more careful and potentially expensive hyperparameter search. Whereas mini-batch training is consistently less sensitive to changes of those parameters than full-graph training, across all the GNN models we consider. In addition, mini-batch training is much less prone to the oversmoothing problem (Rusch et al., 2023) that affects deep GNN models with many layers.
6. Conclusion
This work is motivated by the need to understand the state of the art in GNN training systems, given the absence of consistent empirical evidence comparing full-graph and mini-batch training systems. It first conducted an in-depth literature survey and investigated the reasons for the inconsistent claims and the lack of empirical evidence in the literature. It further provided a thorough empirical evaluation comparing representative systems supporting the two approaches on various GNNs and datasets.
The study reveals that the mini-batch training systems we considered consistently achieve faster convergence time, measured as time-to-accuracy, compared to full-graph training, regardless of the choice of the dataset, GNN models, and hardware settings. The study also shows that using time-to-accuracy as a metric is more relevant than comparing epoch times for systems that have different convergence criteria. Systems of both types can achieve similar accuracy when hyperparameter tuning is performed separately for full-graph and mini-batch training.
Our work stresses the importance of comparing GNN training systems of both classes in future work in this area. Using time-to-accuracy as the main metric for comparison across areas is important, as is performing separate hyperparameter tuning for the two training methods. These findings are relevant to both machine learning practitioners who utilize GNN training systems and systems researchers involved in their design and evaluation.
Appendix
The results for highest achievable accuracy for GAT and GCN model are shown in Table 11 and Table 12. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
Table 13 shows the best architectures obtained for every model-dataset combination.
Figure 7 and Figure 8 illustrate the best test accuracy when training the GCN and GAT respectively. The top row shows mini-batch training experiments and the bottom row shows synchronous full-graph training.
Table 11. Convergence for GAT model
Pubmed Ogbn-arxiv Reddit Ogbn-products
FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
FG-train $77.70\pm 00.50$ $77.20\pm 00.34$ $71.74\pm 00.22$ $67.77\pm 00.31$ $95.17\pm 00.20$ $93.00\pm 00.23$ $74.43\pm 00.05$ $73.33\pm 00.15$
MB-train $75.80\pm 00.06$ $78.50\pm 00.03$ $70.19\pm 00.16$ $70.63\pm 00.12$ $94.58\pm 00.14$ $95.58\pm 00.13$ $74.27\pm 00.07$ $74.57\pm 00.07$
Table 12. Convergence for GCN model
Pubmed Ogbn-arxiv Reddit Ogbn-products
FG-search MB-search FG-search MB-search FG-search MB-search FG-search MB-search
FG-train $79.40\pm 00.12$ $75.10\pm 00.37$ $71.54\pm 00.21$ $69.32\pm 00.22$ $94.67\pm 00.20$ $93.64\pm 00.10$ $75.71\pm 00.18$ $73.81\pm 00.13$
MB-train $78.20\pm 00.16$ $78.70\pm 00.23$ $70.54\pm 00.16$ $71.00\pm 00.16$ $93.31\pm 00.14$ $94.41\pm 00.14$ $75.27\pm 00.07$ $79.31\pm 00.04$
Table 13. Best architectures
Model Training technique Architecture Pubmed Ogbn-arxiv Reddit Ogbn-products
GraphSAGE FG Layers 4 3 4 5
Hidden size 64 512 1024 512
MB Layers 2 5 4 5
Hidden size 256 128 512 256
GAT FG Layers 3 3 2 2
Hidden size 1024 1024 1024 256
Num heads 1 2 2 2
MB Layers 2 4 2 3
Hidden size 1024 256 512 128
Num heads 4 2 2 2
GCN FG Layers 2 2 2 3
Hidden size 512 1024 1024 512
MB Layers 6 2 2 2
Hidden size 64 1024 512 512
Figure 7. Ablation study with number of layers and hidden size for GCN model. Legend represents the size of each hidden layer. X-axis shows the number of layers. Y-axis shows the test accuracy. Top row is for mini-batch training and bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Figure 8. Ablation study with number of layers and hidden size for GAT model. Legend represents the size of each hidden layer. X-axis shows the number of layers. Y-axis shows the test accuracy. Top row is for mini-batch training and bottom row is for full graph training. We observe that mini-batch training is robust to architectural changes, whereas full-graph training has a lot of variation.
Table 14 shows the results of the time to accuracy experiment for GAT model on a single host with NVLink connectivity.
Table 14. Time to accuracy for GAT on a single NVLink host with 4 GPUs. Architectures (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 6, 25), Reddit - (2, 1024, 4, 20), ogbn-products - (3, 128, 3, 20). ET - Epoch time, TTA - Time to accuracy. For ogbn-papers100m DGL uses UVA mode for sampling and Quiver performs distributed caching.
System ogbn-arxiv reddit ogbn-products
ET (s) PipeGCN (FG, Sync) 0.625 1.801 1.033
PipeGCN (FG, Async) 0.469 1.312 0.614
BNS-GCN (FG, Sync) 0.420 1.395 0.629
DGL (MB, Sync) 3.871 6.510 2.730
Quiver (MB, Sync) 3.596 6.779 2.974
TTA (s) PipeGCN (FG, Sync) 275.04 1728.58 413.28
PipeGCN (FG, Async) 211.14 1312.15 233.24
BNS-GCN (FG, Sync) 151.20 1395.40 132.13
DGL (MB, Sync) 42.58 97.65 40.95
Quiver (MB, Sync) 35.96 101.68 44.61
Table 15 shows the results for the scalability experiments for GAT model on PCIe setup and Table 16 shows the results of the scalability experiment on the NVLink setup.
Table 15. Single-host scalability with the GAT model using the PCIe host. Times in seconds. (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 2, 25), Reddit - (2, 1024, 1, 15), ogbn-products - (3, 128, 1, 10), Orkut - (2, 128, 1, 10). ET - Epoch time, TTA - Time To Accuracy.
Dataset ogbn-arxiv Reddit ogbn-products Orkut
System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET
PipeGCN (Sync) 1 2.225 979.09 1.0x 5.857 5622.53 1.0x 3.510 1404.12 1.0x OOM
2 1.514 665.98 1.5x 4.035 3873.98 1.5x 3.381 1352.32 1.0x 2.726
3 1.323 581.98 1.7x 4.034 3872.83 1.5x 2.748 1099.16 1.3x 2.425
4 1.072 471.55 2.1x 2.890 2774.78 2.0x 2.452 980.60 1.4x 2.323
PipeGCN (Async) 1 2.218 998.01 1.0x 5.856 5856.10 1.0x 3.503 1331.10 1.0x OOM
2 1.334 600.08 1.7x 3.269 3269.20 1.8x 2.762 1049.37 1.3x 2.479
3 1.026 461.79 2.2x 3.109 3109.20 1.9x 1.972 749.17 1.8x 1.898
4 0.805 362.03 2.8x 1.879 1879.10 3.1x 1.523 578.78 2.3x 1.499
BNS-GCN (Sync) 1 2.337 841.43 1.0x 6.232 6231.60 1.0x 4.010 842.10 1.0x OOM
2 1.345 484.02 1.7x 3.472 3472.20 1.8x 3.466 727.82 1.2x 2.860
3 0.980 352.76 2.4x 2.697 2697.10 2.3x 2.366 496.92 1.7x 2.278
4 0.778 280.08 3.0x 2.193 2192.70 2.8x 1.710 359.06 2.3x 1.583
DGL (Sync) 1 15.401 154.01 1.0x 17.468 262.01 1.0x 18.394 367.88 1.0x 22.465
2 8.888 88.88 1.7x 8.575 128.62 2.0x 9.328 186.57 2.0x 11.580
3 6.146 61.46 2.5x 5.722 85.83 3.1x 6.246 131.17 2.8x 7.544
4 4.627 46.27 3.3x 4.594 68.91 3.8x 4.715 99.02 3.7x 6.444
Quiver (Sync) 1 16.230 162.30 1.0x 18.914 283.71 1.0x 17.457 349.15 1.0x 24.222
2 8.286 91.15 1.8x 9.002 135.03 2.1x 8.783 175.66 2.0x 14.134
3 5.737 68.85 2.4x 6.972 111.55 2.5x 5.633 118.29 3.0x 9.048
4 4.256 51.08 3.2x 4.595 87.31 3.2x 4.695 98.59 3.5x 6.695
Table 16. Single-host scalability with the GraphSAGE model using the NVLink host. Times in seconds. (#Layers, hidden size, fanout): ogbn-arxiv - (3, 1024, 25), Reddit - (2, 1024, 20), ogbn-products - (3, 128, 20), Orkut - (3, 128, 5). ET - Epoch time, TTA - Time to Accuracy.
Dataset ogbn-arxiv Reddit ogbn-products Orkut
System # of GPUs ET TTA TTA Speed-up ET TTA TTA Speed-up ET TTA TTA Speed-up ET
PipeGCN (Sync) 1 0.181 39.90 1.0x 0.433 121.16 1.0x 0.270 107.90 1.0x 0.466
2 0.182 40.04 1.0x 0.746 208.78 0.6x 0.398 159.28 0.7x 0.452
3 0.315 69.24 0.6x 0.735 205.68 0.6x 0.306 122.34 0.9x 1.312
4 0.211 46.51 0.9x 0.776 217.40 0.6x 0.518 207.20 0.5x 0.995
PipeGCN (Async) 1 0.182 56.28 1.0x 0.432 120.94 1.0x 0.270 102.48 1.0x 0.467
2 0.098 30.38 1.9x 0.255 71.41 1.7x 0.164 62.24 1.6x 0.281
3 0.071 21.97 2.6x 0.212 59.37 2.0x 0.114 43.17 2.4x 0.234
4 0.057 17.53 3.2x 0.130 36.42 3.3x 0.089 33.71 3.0x 0.183
BNS-GCN (Sync) 1 0.188 50.74 1.0x 0.369 58.97 1.0x 0.363 76.31 1.0x 0.536
2 0.106 28.71 1.8x 0.227 36.29 1.6x 0.219 46.03 1.7x 0.322
3 0.083 22.45 2.3x 0.183 29.33 2.0x 0.162 34.00 2.2x 0.289
4 0.069 18.56 2.7x 0.157 25.07 2.4x 0.136 28.64 2.7x 0.225
DGL (Sync) 1 1.519 7.60 1.0x 1.571 7.85 1.0x 2.819 42.28 1.0x 7.182
2 0.842 4.21 1.8x 0.846 4.23 1.9x 1.389 20.84 2.0x 3.606
3 0.577 2.89 2.6x 0.576 2.88 2.7x 0.926 13.90 3.0x 2.390
4 0.457 2.28 3.3x 0.490 2.45 3.2x 0.699 10.49 4.0x 1.842
Quiver (Sync) 1 0.906 9.06 1.0x 2.478 12.39 1.0x 2.463 36.94 1.0x 6.301
2 0.481 4.81 1.9x 1.278 6.39 1.9x 1.250 18.75 2.0x 3.317
3 0.333 3.33 2.7x 0.901 4.51 2.8x 0.861 12.92 2.9x 2.279
4 0.260 2.60 3.5x 0.731 3.66 3.4x 0.643 9.65 3.8x 1.743